OpenStudy (anonymous):

Which numbers between 40 and 49 have 2 and 3 in their prime factorization?

jimthompson5910 (jim_thompson5910):

Any number that is even will have 2 in its prime factorization. For instance, 46 is even and 46 = 2*23 Any number that has its digits add up to 3 (or any multiple of 3) will have 3 in its prime factorization. For instance, the digits of 39 add to 3+9 = 12 which is a multiple of 3 (also notice that the digits of 12 do the same 1+2 = 3)

OpenStudy (anonymous):

42,48

OpenStudy (anonymous):

ok, so any number 2 or 3 can divide equally into?

jimthompson5910 (jim_thompson5910):

yes, or put another way, any number that 6 is a factor of (since 2*3 = 6). In this case, 42 = 6*7 and 48 = 6*8 and these are the only two numbers in this interval that meet these requirements.